### v1

@@ Appendix H Theoretical analysis of the expressivity of our QRC for time-independent signals
The ability of the QRC to perform better than an optimal linear layer on the input is in the reservoir’s ability to express many non-linear functions of the input—its expressivity. Here, we quantitatively characterize the class of functions which can be represented by the oscillator component of the QRC for a time-independent input. In this regime, the input can be represented by two variables: the values of the in-phase and quadrature components. The output feature vector from the QRC is then a function of these two variables.
@@
where we have written $\alpha=|\alpha|e^{i\phi_{\alpha}}$, $\beta=|\beta|e^{i\phi_{\beta}}$, and set $|\alpha|=1/2$. Choosing different values of $\phi_{\alpha}$ gives rise to different output features of the QRC. In this experiment, we pick $\phi_{\alpha}\in\{0,\pi/2\}$, but in principle, one can add to the feature vector with more choices of $\phi_{\alpha}$. For example, one can choose $\phi_{\alpha}\in\{0,\omega,2\omega,\dots,(r-1)\omega\}$ where $\omega=\frac{2\pi}{r}$. The final output after the linear layer is an arbitrary linear combination of all the $p_{\alpha}(\beta)$ functions.
@@
Intuitively, the larger $r$, the more expressive the function space spanned by these features. Furthermore, the higher order central moments allow the output feature vector to represent powers of this probability: $p_{\alpha}(\beta)^{n}$, for moments upto the nth order. We have shown that the qubit measurements extract the phase information of the input complex number $\beta$. Below we will focus on the oscillator parity measurement which is sensitive to the magnitude of $\beta$. Recall that the post-measurement (unnormalized) state of the cavity can be described by a sequence of alternating displacements and parity measurements (Eq. 31):
@@
Appendix I Classicalized reservoir with the Maxwell–Bloch approximation
To investigate the role that quantumness plays in our reservoir’s performance, we performed a simulation of our quantum reservoir in a way that its behavior is classicalized, using the Maxwell–Bloch approximation [5]. This is a mean-field approximation that removes entanglement between qubit and cavity and has the cavity state remain coherent. The statistics extracted from the simulation are the same—namely probability of qubit excitation and cavity parity—but the underlying measurements are assumed to have no effect on the qubit or cavity. That is, the classicalized version of the reservoir removes the effects of measurement backaction as well.
Setup: Hamiltonian and Lindbladian
The Lindblad master equation [lindbladmasterequation] for the Hamiltonian 1 can be written, then as:
$\dot{\rho}=-2\pi i[H,\rho]+\frac{1}{T_{1,\text{storage}}}D_{\rho}[a]+\frac{1}{T_{\phi,\text{qubit}}}D_{\rho}[\sigma_{z}]+\frac{1}{T_{1,\text{qubit}}}\left((1-\bar{n})D_{\rho}[\sigma]+\bar{n}D_{\rho}[\sigma^{\dagger}]\right)$ (51)
$D_{\rho}[L]=L\rho L^{\dagger}-\frac{1}{2}\left(L^{\dagger}L\rho+\rho L^{\dagger}L\right)$ (52)
$\frac{1}{T_{\phi,\text{qubit}}}=\frac{1}{T_{2,\text{qubit}}}-\frac{1}{2*T_{1,\text{qubit}}}$ (53)
Here, $\bar{n}$ is the expected thermal occupation of the qubit. This quantity, along with $T_{1}$,$T_{2}$ of the storage and qubit are selected to match those of the experiment. The $T_{2}$ and thermal occupation of the cavity are negligible compared to all other terms in the time scale of the experiment and so are neglected.
For an operator whose expected value we track, we can take an equation of motion with respect to that expected value:
$\dot{\langle O\rangle}=-2\pi i\langle[O,H]\rangle+\frac{1}{T_{1,\text{storage}}}\langle\tilde{D}_{O}[a]\rangle+\frac{1}{T_{\phi,\text{qubit}}}\langle\tilde{D}_{O}[\sigma_{z}]\rangle+\frac{1}{T_{1,\text{qubit}}}\left((1-\bar{n})\langle\tilde{D}_{O}[\sigma]\rangle+\bar{n}\langle\tilde{D}_{O}[\sigma^{\dagger}]\rangle\right)$ (54)
$\tilde{D}_{O}[L]=L^{\dagger}OL-\frac{1}{2}\left(L^{\dagger}LO+OL^{\dagger}L\right)$ (55)
Equations of Motion
Tracking only first-order central moments (the means), we only need to keep track of $\langle a\rangle$,$\langle\sigma\rangle$, and $\langle\sigma_{z}\rangle$. We assume that there is no entanglement between the qubit and cavity - meaning that at all times for any operator $Q$ on the qubit and $C$ on the cavity we have $\langle QC\rangle=\langle Q\rangle\langle C\rangle$. We also assume that the state in the cavity remains coherent and therefore $\langle a^{\dagger}a\rangle=|\langle a\rangle|^{2}$. Using Eq. 54 we can derive the following equations of motion:
$\dot{\langle a\rangle}=2\pi\times\left(-i\frac{\chi}{2}\langle\sigma_{z}\rangle\langle a\rangle-id_{\text{storage}}^{*}(t)\right)-\frac{1}{2T_{1,\text{storage}}}\langle a\rangle$ (56)
$\dot{\langle\sigma\rangle}=2\pi\times\left(id_{\text{qubit}}^{*}(t)\langle\sigma_{z}\rangle-i\chi\langle\sigma\rangle|\langle a\rangle|^{2}\right)-\frac{2}{T_{\phi,\text{qubit}}}\langle\sigma\rangle-\frac{1}{2T_{1,\text{qubit}}}\langle\sigma\rangle$ (57)
$\dot{\langle\sigma_{z}\rangle}=2\pi\times\left(2i(d_{\text{qubit}}(t)\langle\sigma\rangle-d_{\text{qubit}}^{*}(t)\langle\sigma\rangle^{*})\right)-\frac{1}{T_{1,\text{qubit}}}\left(\langle\sigma_{z}\rangle-(2\bar{n}-1)\right)$ (58)
Gate effects
For operations in our reservoir which take place over very small time scales compared to the period of the cross-Kerr interaction term, such as qubit gates, we implement these as gates directly in our system.
For a $\pi/2$-pulse along the $x$-axis, the resulting unitary is given by $\frac{1}{\sqrt{2}}(I-i\sigma_{j})$. Within all expectations that include them, this results in the transformation:
$\frac{1}{2}\left(\sigma+i[\sigma_{j},\sigma]+\sigma_{j}\sigma\sigma_{j}\right)$
From this, we get, for $j=x$,
$\begin{split}\sigma\rightarrow\frac{\sigma+\sigma^{\dagger}}{2}-i\frac{\sigma_{z}}{2},\ \ \ \ \sigma_{z}\rightarrow\sigma_{y}=i\sigma-i\sigma^{\dagger}\end{split}$ (59)
and for $j=y$,
$\begin{split}\sigma\rightarrow-\frac{\sigma_{z}}{2}+\frac{\sigma-\sigma^{\dagger}}{2},\ \ \ \ \sigma_{z}\rightarrow-\sigma_{x}=-\sigma-\sigma^{\dagger}\end{split}$ (60)
Finally, for the simple $X_{\pi}$ pulse, the transformation is simply:
$\begin{split}\sigma\rightarrow\sigma^{\dagger},\sigma_{z}\rightarrow-\sigma_{z}\end{split}$ (61)
Figure M19: Performance of our reservoir protocol when the underlying system is governed by the Maxwell-Bloch equations of motion given by Eqs. 56,57,58, with the features used being third-order central moments between the outcome of four measurement pairs (cavity parity and qubit excitation). The subplots show performance for (a) spiral classification and (b) radio-frequency communication modulation protocol classification.
Finally, as for the CNOD operation, the way this was implemented in our experimental reservoir involved two opposite displacements at the same frequency, with a $\pi$-pulse on the qubit in between. Note, however, that in the mean field approximation, $\langle\sigma_{z}a\rangle=\langle\sigma_{z}\rangle\langle a\rangle$, and since our qubit was on the Bloch sphere equator during exposure to data and application of CNOD, this term goes to zero (and any deviation from the Bloch equator is due to noise and contains no information). Therefore, $\langle a\rangle$ depends on the drive term only, and CNOD has no effect on the cavity, as the cavity is displaced and then undisplaced by an equal amount. Since the only significant term affecting the qubit on the time-scale of the CNOD is the $|\langle a\rangle|^{2}$ term, which increases and decreases symmetrically about the $\pi$ pulse applied to the qubit, the effect of the cavity is cancelled out and the entire procedure is equivalent to only a $\pi$-pulse on the qubit. This is, therefore, the only part our CNOD that survived in the “classicalized” version of our reservoir control protocol.
Transferring our reservoir controls into the Maxwell-Bloch setting: investigating our protocol’s performance in a classical setting
As seen in Fig. M19(a), simulations of our reservoir using Maxwell-Bloch for classifying spiral, yielded classification accuracies that were no better than that of random guessing. The reason for this is that every point in the spiral has a point in the opposite class with identical magnitude but opposite phase in the quadrature space. However, the Maxwell-Bloch version of our reservoir is unable to differentiate two time-independent signals with identical magnitudes. As discussed in the previous section, our version of CNOD only affects the qubit with a $\pi$-pulse operation. Therefore, the only changes to the cavity are due to displacement caused by the data signal as a driving term, as seen from Eq. 56 since $\langle\sigma_{z}\rangle=0$ and so the qubit state does not influence our cavity during this stage. Consequently, when we measure cavity parity, this contains information about the magnitude of the signal but not its quadrature phase.
Depth, Reservoir size (R) $w_{in}$ sparsity $\rho$ $\gamma$ a
depth=4, $R\geq 64$ 1.0 1.0 1.0 1.0 0
depth=4, $R=32$ 1.0 0.2 1.0 1.0 0
depth=4, $R=16$ 1.0 0.6 1.0 0.9 0.2
depth=4, $R=8$ 1.0 0.6 0.9 0.8 0.5
depth=3, $R\geq 64$ 1.0 0.2 1.0 1.0 0.5
depth=3, $R=32$ 1.0 0.3 1.0 1.0 0
depth=3, $R=16$ 1.0 0.45 1.0 1.0 0.2
depth=3, $R=8$ 1.0 1.0 1.0 1.0 0.0
depth=2, $R\geq 64$ 1.0 0.2 1.0 1.0 0.5
depth=2, $R=32$ 1.0 0.3 1.0 1.0 0
depth=2, $R=16$ 1.0 0.45 1.0 1.0 0.2
depth=2, $R=8$ 1.0 1.0 1.0 1.0 0.0
Table S2: Hyper-parameters that were empirically found to give the best average spiral classification accuracy for each depth and size of reservoir.
The qubit, with the drives turned off, only has a dependence on the cavity in terms of $\langle\sigma_{z}\rangle|\langle a\rangle|^{2}$ as seen in Eq. 58. The qubit excitation probability, which we will obtain later after a $Y_{\frac{\pi}{2}}$ pulse is a nonlinear function of the integral of the cavity’s expected photon number over time. Therefore, information about the signal quadrature phase is absent from qubit measurements as well.
Importantly, this is a flaw that our particular reservoir protocol suffers from in the Maxwell-Bloch setting. For instance, having unconditional displacements, that differ for each run of the reservoir, could help turn quadrature phase information in the signal into information about the storage cavity’s displacement.
This limitation interestingly had a very minimal effect on the performance of the classicalized reservoir when it came to classifying radio-frequency communication modulations. This can be seen in Fig. M19(b), whose shot-limited accuracies were close to those of the quantum reservoir simulation accuracies of Fig. M15(b) with the same number of central moments. This is likely because differences between consecutive quadrature phases of incoming signals now introduce interactions that affect both qubit and cavity.
Another point noticed in our investigation with Maxwell-Bloch was that, while the performance hierarchy of third-order-cumulant-based features outperforming the sampling feature was maintained (see Fig. M13), the advantage was dramatically lower. The lack of variation is the reason why only third-order central moments are plotted in Fig. M19, as the performance curves would lie on top of each other. There are several reasons for this. Firstly, there are no longer higher-order correlations present due to the effects of entanglement between excitation measurements of the qubit and parity measurements of the cavity. Secondly, since this framework does not consider the effects of measurement back-action this removes correlation structures across measurement times that the quantum version takes advantage of. Nevertheless, due to some degree of correlation across measurements due to the signal itself, a weak advantage to considering higher-order central moments, and limitations of shot noise, did result in advantages in classification accuracy. This highlights the computational benefits our system reaps from considering signal correlations and central moments emerge from the information contained there, rather than a consequence of introducing the nonlinearity required to compute these values (since introducing the same correlation features for Maxwell-Bloch did not substantially change accuracy).
Appendix J Leaky Echo State Networks (LESN)
@@ Appendix I Classicalized reservoir with the Maxwell–Bloch approximation
Leaky echo state networks [JAEGER2007] are a generalization of echo state networks (ESN) [jaeger2001echo] that were found to outperform their parent design in prediction and classification of slow dynamic systems, noisy time series and time-warped dynamic patterns [LESNREVIEW2022]. Given a sequence of inputs $\{u_{n}\}_{n=1}^{N},u_{n}\in\mathbb{R}^{D}$, the state of the LESN reservoir after the $n^{\text{th}}$, $x_{n}$, is given by the following equation:
@@
Here, $a,\gamma$ are fixed hyper-parameters in $[0,1]$, and $f$ is a nonlinear activation function. $W_{\text{in}}$ is the $D\times R$ “encoding” matrix whose elements are selected uniformly at random from the interval $[-w_{in},w_{in}]$, where $D$ is the dimension of the input, $R$ is the dimension of the reservoir, and $w_{in}$ is a fixed hyper-parameter. $W_{\text{res}}$ is the $R\times R$ “reservoir” matrix. This matrix is constructed by first generating a matrix $W_{R}$, which is a random matrix whose elements are chosen to be zero with probability $1-p_{s}$ and a number sampled uniformly from the interval $[-1,1]$ with probability $p_{s}$. The largest-magnitude singular value of this matrix, $\lambda_{\text{max}}(W_{R})$ is computed, and the reservoir matrix $W_{\text{res}}$ is defined as:
@@
where $W_{\text{train}}$ is a $R\times C$ trainable linear layer, where $C$ is the dimension of the desired output vector.
@@
Figure M20: Mean spiral classification accuracies and their standard deviations over 100 randomly generated LESN’s. Hyper-parameters for all reservoir sizes $R$ (8,16,32,64,128,256) and depths (2,3,4) are specified in Tab. S2.

### v2

@@ Appendix G Theoretical analysis of the expressivity of our QRC for time-independent signals
The ability of the QRC to perform better than an optimal linear layer on the input lies in the reservoir’s ability to express many non-linear functions of the input—its expressivity. Here, we quantitatively characterize the class of functions which can be represented by the oscillator component of the QRC for a time-independent input. In this regime, the input can be represented by two variables: the values of the in-phase and quadrature components. The output feature vector from the QRC is then a function of these two variables.
@@
We denote $\alpha=|\alpha|e^{i\phi_{\alpha}}$, $\beta=|\beta|e^{i\phi_{\beta}}$, and set $|\alpha|=1/2$. Choosing different values of $\phi_{\alpha}$ gives rise to different output features of the QRC. In this experiment, we pick $\phi_{\alpha}\in\{0,\pi/2\}$, but in principle, one can add to the feature vector with more choices of $\phi_{\alpha}$. For example, one can choose $\phi_{\alpha}\in\{0,\omega,2\omega,\dots,(r-1)\omega\}$ where $\omega=\frac{2\pi}{r}$. The final output after the linear layer is an arbitrary linear combination of all the $p_{\alpha}(\beta)$ functions.
@@
Intuitively, the larger $r$, the more expressive the function space spanned by these features. Furthermore, the higher-order central moments allow the output feature vector to represent powers of this probability: $p_{\alpha}(\beta)^{n}$, for moments up-to the nth-order. We have shown that the qubit measurements extract the phase information of the input complex number $\beta$. Below we will focus on the oscillator parity measurement which is sensitive to the magnitude of $\beta$. Recall that the post-measurement (unnormalized) state of the cavity can be described by a sequence of alternating displacements and parity measurements (Eq. 31):
@@
Appendix H Leaky Echo State Networks (LESN)
@@
Leaky echo state networks [JAEGER2007] are a generalization of echo state networks (ESN) [jaeger2001echo] that were found to outperform their parent design in prediction and classification of slow dynamic systems, noisy time series and time-warped dynamic patterns [LESNREVIEW2022]. Given a sequence of inputs $\{u_{n}\}_{n=1}^{N},u_{n}\in\mathbb{R}^{D}$, the state of the LESN reservoir after the $n^{\text{th}}$ input $u_{n}$, $x_{n}$, is given by the following equation:
@@
Here, $a,\gamma$ are fixed hyper-parameters in $[0,1]$, and $f$ is a nonlinear activation function. $W_{\text{in}}$ is the $R\times D$ “encoding” matrix whose elements are selected uniformly at random from the interval $[-w_{in},w_{in}]$, where $D$ is the dimension of the input, $R$ is the dimension of the reservoir, and $w_{in}$ is a fixed hyper-parameter. $W_{\text{res}}$ is the $R\times R$ “reservoir” matrix. This matrix is constructed by first generating a matrix $W_{R}$, which is a random matrix whose elements are chosen to be zero with probability $1-p_{s}$ and a number sampled uniformly from the interval $[-1,1]$ with probability $p_{s}$. The largest-magnitude singular value of this matrix, $\lambda_{\text{max}}(W_{R})$ is computed, and the reservoir matrix $W_{\text{res}}$ is defined as:
@@
where $W_{\text{train}}$ is a $C\times R$ trainable linear layer, where $C$ is the dimension of the desired output vector.
@@
Figure S19: Mean spiral classification accuracies and their standard deviations over 100 randomly generated LESN’s.
@@
In the conventional view of reservoirs, the data must be sent into into a higher dimensional space where linear separability becomes possible [LESNREVIEW2022]. The dimensionality of our Hilbert space, given by the two dimensions of the qubit and the approximately 16 occupied levels of our storage resonator, limits the complex degrees of freedom we have available in encoding our data as the final, measured state. Consequently, we use this total Hilbert space dimension (times two due to complex amplitudes) as a proxy for the quantum resources used in terms of reservoir dimensionality. Given the roughly $2\times 16$ dimensions of Hilbert space used by our reservoir, achieving at least the computational capacity of a 64-dimensional LESN reservoir is on the order of what would be expected for a large shot number. Indeed, the point of this comparison is to demonstrate that the computations performed by our reservoir cannot be trivially replicated by a digital reservoir with fewer resources with the same performance.